A patient takes a pill that holds 50 mg of medicine in it. The pill is metabolized, or used up, at a rate of 55% every hour.
a. Rounded to the nearest hundredth, how much of the pill has been used by the body after four hours?
1
Expert's answer
2020-06-14T16:07:14-0400

So 55 % percent is used in every hour

After 1 hour

"55\ \%" of 50 mg will be used then "\frac{55}{100}\times50=27.5mg"

now we are left with 50-27.5= 22.5 mg


after 2 hours

55 % of 22.5 mg will be used ="\frac{55}{100}\times22.5=12.375 mg"

now we are left with 22.5-12.375=10.125 mg


after 3 hours

we will be left with 4.556 mg


after 4 hours

we will be left with 2.05 mg


total pill consumed = 50-2.05=47.95 mg